OpenStudy (anonymous):

How to write an equation of a line that passes through (4,3) and (-8,6)

OpenStudy (amistre64):

define the slope of the line between the given points, and use it to construct the point slope form of the line.

jigglypuff314 (jigglypuff314):

\[slope = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (amistre64):

y = m(x-xo) + yo

OpenStudy (texaschic101):

first you need to find the slope by using the slope formula slope(m) = (y2 - y1) / (x2 - x1) (4,3) x1 = 4 and y1 = 3 (-8,6) x2 = -8 and y2 = 6 now sub in your info into the formula... m = (y2 - y1) / (x2 - x1) m = (6 - 3) / (-8 - 4) can you continue and find the slope from here ?

OpenStudy (texaschic101):

Once you have the slope, you will use point slope formula... y - y1 = m(x - x1) slope(m) = (whatever answer you got for the slope) and you can use either set of points... (4,3) x1 = 4 and y1 = 3 or (-8,6) x1 = -8 and y1 = 6 pick ONE of those sets of points now just sub your info into the formula. Once that is done, solve for y and that will be your equation. It will be in y = mx + b form.

OpenStudy (anonymous):

thanks you guys!